l he data from 200 machined parts are summarized as follows: Depth of Bore Edge Condition Coarse Moderate Smooth Above Target Below Target 10 23 84 unt 15 oard 46 ses (a) What is the probability that a part selected has a moderate edge condition and a below-target bore depth? Round your answer to two decimal places (eg.98.76). dar b) What is the probability that a part selected has a moderate edge condition or a below-target bore depth? Round your answer to two decimal places (e-g.98.76) (c) What is the probability that a part selected does not have a moderate edge condition or does not have a below-target bore depth? Round your answer to two decimal places (e-g.98.76).
Answer #1

(a)
Probability that a part selected has a moderate edge condition and a below target bore depth
= Frequency of moderate edge condition and a below target bore depth / Total frequency
= 23 / 200 = 0.115 = 0.12 (Rounding to 2 decimal places)

(b)
Probability that a part selected has a moderate edge condition or a below target bore depth
= Frequency of moderate edge condition or a below target bore depth / Total frequency
= (22 + 23 + 10 + 84) / 200 = 0.695 = 0.70 (Rounding to 2 decimal places)

(c)
Probability that a part selected does not have a moderate edge condition or does not have below target bore depth
= Frequency of observations except moderate edge condition or a below target bore depth / Total frequency
= Frequency of coarse/smooth edge condition or above target bore depth / Total frequency
= (15 + 46) / 200 = 0.305 = 0.31 (Rounding to 2 decimal places)